MICROBIO HELP: You want to grow up some bacteria to bring to class to show your friends, and you want to have enough to give everyone in the class 100 bacteria. If you are starting with only 5, how many generation times should you allow so you will have enough? (assume there are 130people in the class). If you knew it took 1 hour and 39 minutes to grow all your bacteria, what is the generation time?

Helpful information provided:

Calculating growth parameters

A.Mathematically

1.N = 2^nNo

N = final number.

n = number of generations

No = starting number of cells

d) Example. If you start with 3 cells, how many cells will you have after 4

generations.

2.Generation time

(also called doubling time)

g = t/n

Where t = time it took and n is the number of generations

Ex/ you started your culture at 9 AM, and 4 hours later the bacteria had

gone through 5 generations. What is the generation time? g = 4/5 = 0.8 hours
Answer #1

Using the formula: N=2n N0

where, N= 130 X 100 =13000 N0 = 5

So, 13000 = 2n  X 5

2n = 13000/5 = 2600

Taking log on both sides,log 2n = log 2600

n log 2 = log 2600

n= log 2600/log 2 = 3.414 /0.301 = 11.342

Number of generations (n) = 11.342
Generation time, g= t/n
t= 1 hour 39 minutes =1.65 hours
g= 1.65/11.342 = 0.145 hours
So the generation time is 0.145 hours.